OpenStudy (anonymous):

Which is greater? Line l has slope > 1. A. Slope of line p B. -1

OpenStudy (anonymous):



Nnesha (nnesha):

line decreasing = negative slope line increasing = positive slope

OpenStudy (anonymous):

Does the perpendicular line have anything to do with this problem?

Nnesha (nnesha):

perpendicular slopes are negative reciprocal

OpenStudy (anonymous):

Hey but the answer says that A is greater than B

imqwerty (imqwerty):

The answer is B -_-

OpenStudy (anonymous):

no it's A

Nnesha (nnesha):

sorry i had to go

imqwerty (imqwerty):

Lets take the slope of line l as n n>1 We knw that the lines r perpendicular so the product of their slopes = - 1 Let the slope of line p be m So n x m= - 1 m= - 1/n Now u knw that n>1 and m= - 1/n What can u cmnt abt the value of m will it be more than - 1 or less than it.